Carla invested $28,000, part at 14% and part at 9%. If the total interest at the end of the year is $3,370, how much did she invest at 14%?

Answer by checkley75(3666)

.14X+.09(28000-X)=3370
.14X+2520-.09X=3370
.05X=3370-2520
.05X=850
X=850/.05
X=17000 INVESTED @ 14%
PROOF
.14*17000+.09*11000=3370
2380+990=3370
3370=3370